# Globular Clusters GMRT Pulsar Search (GCGPS) II: Discovery of five MSPs in M69 and M70

### v1

@@ 4.2 The GC images
The clusters NGC 5986 and NGC 6652 had known MSPs prior to the GCGPS observations. In the case of NGC 5986, the imaging revealed several faint compact sources (with flux densities of a few hundred Jy) located outside the half-light radius, which could be potential MSP candidates. However, the lack of a well-localised position for the previously discovered pulsar in this cluster (J15463747A) did not allow us to associate it with any of the detected sources in the image field. For NGC 6652, similar to the findings of Gautam2022, we detect the source that they associated with J18353259B (marked in the image), but no findings for J18353259A, along with several other faint compact sources that are not associated with any known radio emitters. These additional detections are promising candidates for follow-up studies aimed at identifying new MSPs.
There are now four GCs (NGC 1904, NGC 6093, NGC 6637, and NGC 6681) that host first known MSPs for each GC, discovered through the GCGPS survey. For NGC 1904, the localisation work for the GCGPS-discovered pulsar NGC 1904A (J05242431A) is currently ongoing. We include the image of NGC 1904 in this paper for completeness. In NGC 6093, three MSPs are currently known: the first discovered by the GCGPS survey (PAPER I), and two subsequently found by the TRAPUM survey333See https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html. Of these, only NGC 6093A (J16172258A) has a well-localised position (from PAPER I), which we mark in the image. Similarly, for NGC 6637, the localised MSPs discovered in this work are marked in its image. The image of NGC 6681 is also presented, although the positions of the MSPs discovered by TRAPUM in this cluster (three more after the first three discoveries from the GCGPS survey) are not yet publicly available. Across all four clusters, the images reveal several faint yet compact, unassociated radio sources. These may correspond to the GCGPS-discovered MSPs (for NGC 1904 and NGC 6681), or TRAPUM-discovered MSPs (in NGC 6093 and NGC 6681), whose positions are not yet released, or they could represent entirely new, yet-undiscovered MSPs. These sources warrant targeted follow-up observations.
Except for these six GCs, sixteen additional clusters in the GCGPS survey resulted in non-detections and currently do not host any known MSPs. For most of these clusters, the imaging reveals several previously unassociated compact radio sources, which could plausibly include yet-undiscovered MSPs. To enable dedicated follow-up observations of these clusters with deeper sensitivity, we publish the processed FITS images along with this paper (see DOI: 10.5281/zenodo.17781807, private untill published in ApJ ). We appreciate using these images to identify and prioritise promising targets for future searches.
4.3 Non-detection limits
@@
According to Dewey_1985, for continuum emission, the radiometer equation can be written as:
@@ 5 SUMMARY
The GCGPS observation of the GC M69 resulted in the discovery of two MSPs: M69A is a 3.80 ms MSP (), it was detected with a significant acceleration () during the discovery observation, confirming it’s a binary in a relatively compact orbit; whereas M69B, a 4.80 ms MSP (), was detected with no acceleration, suggesting it is isolated or a binary system in a wide orbit. We localize both MSPs in the image plane (see Figure 2 and Table 2) with arc-sec precision, by simulating the discovery uGMRT beams along with using the simultaneously recorded image data during discovery observation and subsequent follow-up observations.
The observation of M70 led to the discovery of three new millisecond pulsars (MSPs). M70A is a 3.93 ms pulsar with a dispersion measure (DM) of 70.65 pc cm-3. The brightest among them is M70B, a 4.44 ms MSP with a DM of 71.14 pc cm-3, while the faintest is M70C, a 6.06 ms MSP with a DM of 71.06 pc cm-3. None of these sources exhibited detectable acceleration over the 2.5-hour discovery observation, indicating they are either isolated or in binary systems with relatively wide orbits. As in the case of M69, the beam simulation, along with the detection ratios of both beams (CSQ and 3rd arm), indicated that all three MSPs are located within the very central region of M70. Imaging (see Figure 3) revealed a very bright, extended radio source near the center of the globular cluster, which prevented us from confidently associating any of the MSPs with a compact radio source. However, a faint, compact source was also detected almost exactly at the cluster center. This source is a strong candidate for being associated with one of the MSPs. Dedicated timing follow-up is required to determine the precise positions of the MSPs and to confirm any association with the detected radio source.
The imaging of all globular cluster fields observed in the GCGPS survey, including those that did not yield any MSP discoveries (see Figures 5, 6, and 7), revealed several promising targets that could potentially host undiscovered MSPs. These candidates may have been missed in the original GCGPS search due to reduced sensitivity, particularly in the outer regions of the clusters. The processed FITS files corresponding to these images are published with this paper (DOI: 10.5281/zenodo.17781807, private untill published in ApJ) to support more focused, targeted follow-up efforts. Additionally, the imaging data were used to compute non-detection limits for each of the GCGPS-observed globular clusters using the radiometer equation (Equation 3). Based on the respective observational parameters, the non-detection limits (both continuum and beamformed) for each cluster, at the observed uGMRT frequency band, are presented in Table 1.

### v2

@@ IV.2 The GC images
The clusters NGC 5986 and NGC 6652 had known MSPs prior to the GCGPS observations. In the case of NGC 5986, the imaging revealed several faint compact sources (with flux densities of a few hundred Jy) located outside the half-light radius, which could be potential MSP candidates. However, the lack of a well-localised position for the previously discovered pulsar in this cluster (J15463747A) did not allow us to associate it with any of the detected sources in the image field. For NGC 6652, similar to the findings of Gautam et al. (2022), we detect the source that they associated with J18353259B (marked in the image), but no findings for J18353259A, along with several other faint compact sources that are not associated with any known radio emitters. These additional detections are promising candidates for follow-up studies aimed at identifying new MSPs.
There are now four GCs (NGC 1904, NGC 6093, NGC 6637, and NGC 6681) that host first known MSPs for each GC, discovered through the GCGPS survey. For NGC 1904, the localisation work for the GCGPS-discovered pulsar NGC 1904A (J05242431A) is currently ongoing. We include the image of NGC 1904 in this paper for completeness. In NGC 6093, three MSPs are currently known: the first discovered by the GCGPS survey (PAPER I), and two subsequently found by the TRAPUM survey777See https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html. Of these, only NGC 6093A (J16172258A) has a well-localised position (from PAPER I), which we mark in the image. Similarly, for NGC 6637, the localised MSPs discovered in this work are marked in its image. The image of NGC 6681 is also presented, although the positions of the MSPs discovered by TRAPUM in this cluster (three more after the first three discoveries from the GCGPS survey) are not yet publicly available. Across all four clusters, the images reveal several faint yet compact, unassociated radio sources. These may correspond to the GCGPS-discovered MSPs (for NGC 1904 and NGC 6681), or TRAPUM-discovered MSPs (in NGC 6093 and NGC 6681), whose positions are not yet released, or they could represent entirely new, yet-undiscovered MSPs. These sources warrant targeted follow-up observations.
Except for these six GCs, sixteen additional clusters in the GCGPS survey resulted in non-detections and currently do not host any known MSPs. For most of these clusters, the imaging reveals several previously unassociated compact radio sources, which could plausibly include yet-undiscovered MSPs. To enable dedicated follow-up observations of these clusters with deeper sensitivity, we publish the processed FITS images along with this paper (see DOI: 10.5281/zenodo.17781807 ). We appreciate using these images to identify and prioritise promising targets for future searches.
IV.3 Non-detection limits
@@
According to Dewey et al. (1985), for continuum emission, the radiometer equation can be written as:
@@ V SUMMARY
The GCGPS observation of the GC NGC 6637 (M69) resulted in the discovery of two MSPs: M69A is a 3.80 ms MSP (), it was detected with a significant acceleration () during the discovery observation, confirming it’s a binary in a relatively compact orbit; whereas M69B, a 4.80 ms MSP (), was detected with no acceleration, suggesting it is isolated or a binary system in a wide orbit. We localize both MSPs in the image plane (see Figure 2 and Table 2) with arc-sec precision, by simulating the discovery uGMRT beams along with using the simultaneously recorded image data during discovery observation and subsequent follow-up observations.
The observation of NGC 6681 (M70) led to the discovery of three new millisecond pulsars (MSPs). M70A is a 3.93 ms pulsar with a dispersion measure (DM) of 70.65 pc cm-3. The brightest among them is M70B, a 4.44 ms MSP with a DM of 71.14 pc cm-3, while the faintest is M70C, a 6.06 ms MSP with a DM of 71.06 pc cm-3. None of these sources exhibited detectable acceleration over the 2.5-hour discovery observation, indicating they are either isolated or in binary systems with relatively wide orbits. As in the case of M69, the beam simulation, along with the detection ratios of both beams (CSQ and 3rd arm), indicated that all three MSPs are located within the very central region of M70. Imaging (see Figure 3) revealed a very bright, extended radio source near the center of the globular cluster, which prevented us from confidently associating any of the MSPs with a compact radio source. However, a faint, compact source was also detected almost exactly at the cluster center. This source is a strong candidate for being associated with one of the MSPs. Dedicated timing follow-up is required to determine the precise positions of the MSPs and to confirm any association with the detected radio source.
The imaging of all globular cluster fields observed in the GCGPS survey, including those that did not yield any MSP discoveries (see Figures 5, 6, and 7), revealed several promising targets that could potentially host undiscovered MSPs. These candidates may have been missed in the original GCGPS search due to reduced sensitivity, particularly in the outer regions of the clusters. The processed FITS files corresponding to these images are published with this paper (DOI: 10.5281/zenodo.17781807) to support more focused, targeted follow-up efforts. Additionally, the imaging data were used to compute non-detection limits for each of the GCGPS-observed globular clusters using the radiometer equation (Equation 3). Based on the respective observational parameters, the non-detection limits (both continuum and beamformed) for each cluster, at the observed uGMRT frequency band, are presented in Table 1.